please help me solve this problem asap
The first of three consecutive numbers is 15. Write the other two.
ps. I tried my best to find the right topic so im sorry if its wrong

Answer by askmemath(368)

1,2,3 can be called consecutive numbers since each number differs from the previous number by 1
Thus consecutive numbers are X,X+1,X+2
So if 15 is the first of 3 consecutive number the next number should be 15+1 = 16 and the number after that should be 15+2 = 16+1 = 17